OpenStudy (anonymous):

HELP 5 questions! PLEASE ANYONE! WILL FAN AND MEDEL AND RIGHT A GOOD TESTIMONIAL

OpenStudy (anonymous):

X-2 over 7 = 3

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@chase17

OpenStudy (anonymous):

i think the answer is 3 and 2 over 7

OpenStudy (anonymous):

\[3- \left(\begin{matrix}2\\ 7\end{matrix}\right)\]

OpenStudy (anonymous):

=3

OpenStudy (anonymous):

@iGreen